if 2 * 1 = 81 , 3 * 2 = 278 , 2 * 5 = 8125, then 1*3 = ?  
  1. 127
  2. 271
  3. 126
  4. 129
  5. 201

Grade:10

2 Answers

Nishant Vora IIT Patna
9 years ago
2 * 1 = 2^3 * 1^3= 81
3 * 2 =3^3 2^3 =278
2 * 5 = 2^3 5^3= 8125
then 1*3 = 1^3 3^3 = 127

Garv Garg
6 years ago
2×1=813×2=2782×5=81253×1=?We can do cube of all no2³×1³3³×2³2³×5³1³×3³So now we can write812788125 Your answer is 127